OpenStudy (anonymous):

3 2^x = 6 5^x solve for x

OpenStudy (cwrw238):

simplify first by dividing both sides by 3: 2^x =2* 5^x now taks logs: ln 2^x = ln 2 + ln 5^x by laws of logs x ln 2 = ln 2 + x ln 5 can you continue from here

OpenStudy (cwrw238):

just think of ln 2 and ln 5 as numbers (which they are) so - ln 2 =( ln 5 - ln 2) x x = - ln 2 / (ln 5 - ln 2)

OpenStudy (anonymous):

(ln6-ln3)/(ln2-ln5)

OpenStudy (anonymous):

how do you get ln2 on top?

OpenStudy (cwrw238):

right - thats basically what i have

OpenStudy (anonymous):

hows does ln6-ln3 = ln2

OpenStudy (cwrw238):

i'm dividing both sides by ( ln5 - ln2)

OpenStudy (cwrw238):

ln6 - ln3 = ln 6/3

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

yw